OpenStudy (anonymous):

Two complementary angles have measures of 4x + 14 and 3x - 22. What is the measure of the smaller angle?

OpenStudy (lynfran):

complementary angles adds up 90 ...so make an equation and solve for x

OpenStudy (lynfran):

4x+14=90...

OpenStudy (anonymous):

19

OpenStudy (lynfran):

4x+14=90 4x=90-14 4x=76 x=19... do the same to the other and solve for x...

OpenStudy (anonymous):

37.33

OpenStudy (lynfran):

correct now which angle is smaller?

OpenStudy (anonymous):

. 14 b. 70 c. 20 d. 90

OpenStudy (anonymous):

those are my answer options